OpenStudy (mukul123):

How can the equivalent mass of an element be defined with respect to Oxygen? Example???

OpenStudy (aaronq):

setting this proposed mass unit to have the weight of 16 amu 16 amu = 1 "oxygen" unit Eg. Si is 28.8 amu, \(\sf \dfrac{1"oxygen" unit}{16 ~amu}=\dfrac{Si~in "oxygen"units}{28.8~amu}\) \(\sf Si~in "oxygen"units=\dfrac{1"oxygen" unit*28.8~amu}{16 ~amu}=1.8"oxygen" units\)

OpenStudy (welshfella):

I believe that the present equivalent masses of the elements are now based on C 12 isotope not oxygen.

OpenStudy (aaronq):

i think the question is posing a hypothetical scenario